March 2004

Sunday Monday Tuesday Wednesday Thursday Friday Saturday

Survey crew in Hall setting up Network
Survey crew in Hall setting up Network

Electrical crew for Detector Hall rack power installation are hired - spend the day in Training.

T&M crew from Tgt area place steel "road plate" over sump covers at the base of the Shaft.  What's the Story with this Steel Plate?  
Survey crew in Hall setting points with laser tracker
Projected Work:
• Achieve BO
• Start Hall electrical work
• Complete the setup of the Hall survey Network
• Start moving in tools and materials
Electricians start work in Hall.  Continue all week.
Survey crew sights down Hall ventilation shaft to tie-in Hall Network to surface.
Official BO Date.
190 or so punch-list   items left for RBI to do
First day of FNAL ownership - 36 people stand in line for 24 access badges.

Survey crew gets Shaft in AM to finish Hall Network tie-in.

Various installation materials delivered.
Punch-List work by RBI continues.

Delivery of various installation tools & materials, to MSB and on to underground.

Controls & LAN racks delivered to Hall.

We have working phones in MSB, at base of shaft, and in Hall.

Projected Work:
• Install DK endcap
• Install Absorber skids underground
• Start pipefitters on Absorber area piping
• Continue electrical work in Detector Hall

ALL THIS WEEK is an accelerator shutdown - some of the MINOS Inst folks will be spending a day or two at CDF and at Booster.
DK Endcap taken underground, up the hill to Ab Encl, held in place with the big forktruck, and 1st pass welds done.

Problem with big forktruck - it's not happy going up the hill.  Mechanic arrives and works on it. Fixed, but won't be used until paperwork from manufacturer comes in.

Skids delivered to Ab Encl using smaller forktruck.

Electricians in Hall continue with rack AC power.

RBI continues with Punch List work.
DK Endcap welding completed.

2 of 4 Ab skids set in place. Pipefitters hired, get training, and move their tools & materials into Ab Encl.

Electricians continue in Detector Hall. Most conduit and outlet frames in place - wiring for lower-level dnst end racks is pulled.

C. Bower, U. Indiana, measures background rates in Hall, with assistance from E. Mislivec of U Minn Duluth.

A. Erwin and Costas  of U. Wis deliver the Muon detector support frames to MSB.

RBI continues Punch List work.
4 of 4 Ab skids set in place.  Pipefitters start their work.

Big Fork Truck gets a test hauling a 10-ton block up the hill, in the presence of a manufacturer's representative and mechanic. While the load is carried up the hill, they agree that the forktruck is not operating to specs, and it will be sent out for service.

Electricians continue work in Hall.

Hardstand for Ab shield blocks set up on the road leading into the site.

RBI continues Punch List work.
Begin Staging Ab steel Blu Blks to MSB hardstand.

Number of Access Badges raised to 26.   Why 2 more Badges?

Additional steel plate delivered underground for sump area, to fill in the footprint of the Plane Cart. What's the Story with this Steel Plate?

Electricians continue in Hall.

Pipefitters continue in Ab Encl.
Continue staging Blu Blks.

Additional steel plate laid out  around shaft base.

Muon Stand parts moved underground and stored in most dnst Alcove.

Detector support structure work:
• Bookend modified to make shimming for 1st plane easier.
• Stainless steel top rails for east-side laid in place (but not welded yet).

Pipefitters continue in Ab Encl.

Electricians continue in Hall. AC power for lower level Spectrometer racks is complete.
3 PPD techs work on overtime to remove pipes from underground. What are these pipes?
Projected Work:
• Deliver Cart to the Hall (Tuesday)
• Complete tasks required to begin plane installation - setting top rails and shimming bookend
• Repair sump cover supports. What's this repair work?
• Start Near Hall LAN wiring contract.
• Continue pipefitters in Absorber
• Continue electricians in Hall

East side detector support top rail welded into place.  This is the one that doesn't need precise alignment.

Survey crew starts - sets up equipment and establishes the reference line.

Big Fork Truck brought up, loaded onto a truck, and sent back to manufacturer.

LAN wiring contractors start work in the Hall.

Electricians continue in Hall.

Pipefitters continue in Ab Encl.

RBI continues Punch List work.
Detector cart delivered underground (whew!).

Bookend survey measurements completed.

Electricians continue in Hall.

LAN inst. continues.

Pipefitters continue in Ab Encl.

RBI continues Punch List work.
Sump cover support repair work begins.  No shaft crane usage for the remainder of the week.

Surveyors worked with welder to attach stainess top rails. Called away to CDF at 11:30. Back tomorrow.

Tested procedure for placing racks onto upper platform.

LAN wiring installation complete.

Electricians continue in Hall.  All lower level rack outlets completed.

Pipefitters continue in Ab Encl.

Two of the electricians from the Hall work on Ab skid electrical connections.

Continue staging Blu Blks outside.
Sump cover repair work continues.

Survey of west stainless top rail completed, tack welded to hold in place.

Electricians continue in Hall.

Pipefitters continue in Ab. Encl. Vacuum piping nearly complete - waiting on a part. Started on water piping.

Ab Skid electrical work continues.

Electricians install conduit from Ab panelboards to the area where the Muon Detector racks will sit.

Laborers installing a metal dam along the west side Ab Tunnel floor, to prevent water from sheeting across the floor.

Sump cover repair work continues.

West stainless top rail welding complete. Bookend shims welded into place.

Electricians continue in Hall.

Pipefitters continue in Ab Encl.

Ab Skid electrical work complete.

Laborers continue with water dam in Ab Access Tunnel.

All steel Blu Blocks for Absorber are staged on the hardstand outside MSB.

Projected Work:
• Install first Detector plane(s)
• LAN Fiber Termination
• Deliver MINOS LCW skid to Hall
• Complete Rack electrical work in Hall, continue electricians on LCW skid and other  finish-up work
• Continue pipefitters in Absorber area
Sump cover repair work completed.

MINOS LCW skids delivered underground and set into place.

Survey crew in Hall checking top rails and bookend shims now that all are in place. Also set up the measurement system to be used for coil collar alignment during plane installation.

Electricians continue in Hall - DAQ rack power, LCW Skid hookup, and miscellaneous outlets.

Pipefitters continue in Ab Encl.
LAN Fiber termination job begins.

The units for the GPS antenna fiber are installed in racks in MSB and in the Near Hall, as this needs a different type of terminator from the LAN fiber, and it helps the installers to see what the fiber will plug into.

The plane transfer fixture, which is used on the cart underground, is moved underground - it had been sitting on the floor of MSB since last Thurs. (it was not attached to the cart when that went underground).

Electrical work continues.

Pipefitter work continues.
Plane 281 installed.  Moved from NM to MSB in the morning, and moved underground in the afternoon.

LAN fiber termination continues.

Electrical work continues.

Pipefitter work continues.